# CAMA: Enhancing Multimodal In-Context Learning with Context-Aware Modulated Attention

### v2

@@ Abstract
Multimodal in-context learning (ICL) is emerging as a key capability that enables large vision-language models (LVLMs) to adapt to novel tasks without parameter updates, expanding their utility across various real-world applications. However, ICL remains unstable, even with well-matched in-context demonstrations (ICDs), suggesting that LVLMs struggle to fully utilize the provided context. While existing efforts focus on prompt engineering or post‑hoc logit calibration, we instead investigate the underlying attention dynamics to overcome LVLMs’ inherent limitations. We identify two critical deficits in their self-attention that impair effective ICL. To bridge the gap, we propose Context-Aware Modulated Attention (CAMA), a plug-and-play and training-free method that dynamically modulates LVLM’s attention logits based on the input in‑context sequence. CAMA employs a two-stage attention modulation to address both identified deficits, enhancing the focus on semantically significant tokens, particularly visual ones. Across four LVLMs and seven benchmarks, CAMA consistently outperforms vanilla models and baselines, demonstrating great effectiveness and generalization. It can also activate the desired effects of prompt engineering methods and remains robust under diverse sequence configurations. Thus, CAMA paves the way for deeper explorations of attention dynamics to advance multimodal reasoning.
@@ Introduction
Large vision language models (LVLMs) have emerged as powerful tools for multimodal information processing and generation (Zhao et al. 2025). Through large-scale pretraining, they integrate visual and textual signals into the shared representation space of large language models (LLMs) and have achieved notable success across vision-language tasks (Ye et al. 2023; Liu et al. 2023). However, adapting LVLMs to new domains remains challenging due to the high costs of multimodal data preparation and training.
To mitigate these costs, researchers are applying in-context learning (ICL), a technique widely used in LLMs (Brown et al. 2020; Dong et al. 2024), to LVLMs. In ICL, a few in-context demonstrations (ICDs) are incorporated into the input as reference examples, enabling the model to adapt to new tasks by interpreting these demonstrations without parameter updates. Recent advancements in model architectures and training protocols have enabled LVLMs to process multiple images and perform interleaved reasoning, making multimodal ICL practical (Alayrac et al. 2022; Laurençon et al. 2024a; Bai et al. 2025; Chen et al. 2025). These advances expand the practical scope of LVLMs (Doveh et al. 2024; Chen et al. 2024b; Guo et al. 2024).
@@
However, the benefits of multimodal ICL can be diminished by its pronounced instability. Recent studies have found that LVLM ICL performance is highly sensitive to subtle prompt details. Minor changes in the order or formatting of ICDs can cause significant performance swings (Liu et al. 2021; Gao, Fisch, and Chen 2021; Lu et al. 2022). This sensitivity intensifies as the number of ICDs grows or when complex multimodal reasoning is required (Chen et al. 2024d; Li et al. 2025b). To improve multimodal ICL in LVLMs, two main routes have been explored. The first optimizes the prompt. It adds guiding text, highlights key image regions, or selects and orders ICDs with predefined metrics (Li et al. 2023a; Yang et al. 2024b). However, prompt engineering requires substantial prior knowledge, lacks stability, and depends heavily on the base model. The second route edits models’ internal logits, as in contrastive decoding. This approach alters reasoning more directly but needs distorted inputs and extra forward passes for calibration (Liu, Zheng, and Chen 2024; Lee, Tsai, and Chiu 2024). Since practical ICL prioritizes efficiency, our work investigates the following questions: What is the intrinsic limitation underlying the instability of multimodal ICL in LVLMs? Can we address it using a more efficient, training-free method?
In pursuit of these questions, we examine the attention dynamics of LVLMs during multimodal ICL. Because their inputs interleave images and texts, we move beyond the existing focus on single-image scenarios and analyze LVLMs from two distinct perspectives: (1) text-based visual grounding within each image-text pair, and (2) query sample‑driven attention allocation across ICDs. We visualize attention deficits in both aspects through targeted experiments and verify that these deficits lead to insufficient utilization of the in-context sequence by the LVLM, resulting in unstable multimodal ICL. In response to these deficits, we present Context-Aware Modulated Attention (CAMA), a training-free method that dynamically modulates the model’s internal attention logits during inference based on the input context. CAMA addresses the two deficits with a two-stage modulation that targets the shallow and middle layers of the decoder. Stage I applies intra-ICD grounding, highlighting the image tokens in each ICD and in the query sample that best align with their accompanying text, which reduces the attention sink introduced by the interleaved image-text format. Stage II performs query-centric routing, reallocating attention among ICDs in proportion to their value to produce the desired answer, thus improving the use of context. Experiments demonstrate that CAMA delivers consistent performance gains, providing insights for reshaping attention mechanisms to build models with better visual capabilities.
@@
In-context Learning (ICL). ICL enables models to solve unseen tasks by conditioning on an input sequence of input–output examples (i.e., ICDs) without updating any parameters (Brown et al. 2020; Dong et al. 2024). This capability markedly improves their practicality and is especially valued in resource‑intensive multimodal domains (Wies, Levine, and Shashua 2023). Large vision–language models (LVLMs) gain ICL capabilities through targeted pretraining or fine‑tuning on interleaved image–text data, as in Flamingo (Alayrac et al. 2022) and LLaVA-NeXT (Li et al. 2024a). Now, it has been a core ability of commercial models such as QwenVL (Bai et al. 2025) and InternVL (Chen et al. 2025). Consequently, recent work has begun to explore the mechanisms of multimodal ICL (Doveh et al. 2024; Chen et al. 2024b). However, few studies probe internal attention. Cross-modal interactions in multimodal ICL therefore remain only partly understood, which prevents state-of-the-art (SOTA) models from fully exploiting ICL’s potential.
Enhancing Multimodal ICL. Limited use of input‑sequence information by LVLMs is considered a major source of instability in multimodal ICL (Liu et al. 2021; Gao, Fisch, and Chen 2021; Li et al. 2023b). Three solution paths have thus emerged. First, task-oriented datasets together plus instruction tuning (Jiang et al. 2024; Chen et al. 2024c) or direct preference optimization (DPO) (Jia et al. 2025) train LVLMs to reason across multiple ICDs. Second, prompt-level methods optimize ICD selection with metrics such as similarity scores and information entropy (Li et al. 2023a; Yang et al. 2024b; Zhou et al. 2024; Wu et al. 2022), or by training automatic selectors (Li et al. 2025c; Yang et al. 2024a). Third, calibration-based methods adjust LVLMs’ final logits using contrastive decoding (Kim et al. 2024; Fazli, Wei, and Zhu 2025; Li et al. 2025a). While these approaches improve performance, they all face challenges. The first requires extensive curated data and parameter updates, the second is less adaptable to fixed-prompt scenarios, and both depend on the base model. The third needs to design distorted input sequences, plus multiple forward passes. In contrast, our CAMA sidesteps all these drawbacks by modulating attention logits at inference time without additional data or tuning.
@@
which directly reveals both the direction and intensity of interactions between any two tokens during the forward pass. After applying a causal mask followed by softmax, $\mathbf{A}^{l,h}$ becomes the attention matrix that reweights the token values $V^{l,h}$ to produce the head’s attention output.
@@
To understand the instability of multimodal in-context learning in LVLMs, we delve into the model’s attention mechanism to analyze its dynamics in both effective and ineffective scenarios. Our objective is to uncover distinct patterns that distinguish these cases. Specifically, we investigate attention dynamics at two levels: (1) Within each ICD: We examine text-based visual grounding to verify whether the model pays attention to relevant objects in the images. (2) Across ICDs: We evaluate whether the model effectively distributes attention among different ICDs according to their relevance to the query. To ensure generalizable experimental results, we experiment with 3-shot settings on two LVLMs: Llava-NeXT-7B (Liu et al. 2024) and Idefics2-8B (Laurençon et al. 2024b). Query samples and ICDs are taken from the validation and training sets of VQAv2, respectively.
Setups. We pair each query with three ICDs of the same question type (e.g., “How many” or ”Is there”) and process the resulting sequences using both LVLMs. From this pool, we identify 2,500 sequences where LLaVA-NeXT-7B produces a correct answer while Idefics2-8B fails, and another 2,500 sequences with the opposite outcome. These form two distinct sets of 5,000 cases each: a “correct” group (indicating effective ICL) and a “wrong” group (indicating ineffective ICL). For each sequence, we manually annotate bounding boxes on the ICD images based on their corresponding textual descriptions. We then extract attention maps from each model layer, identify the top 20% most-attended regions, and compute their Intersection over Union (IoU) with the annotated boxes to obtain the intra-ICD alignment score $s_{\text{align}}$. Additionally, we sample 10,000 sequences from the original pool and replace two of the three ICDs in each sequence with unrelated images, leaving only one key ICD relevant to the query type. For each of these sequences, we generate three variants by positioning the key ICD in the first, second, and third slots, respectively. Using the same processing pipeline, we construct new correct and wrong groups of 5,000 cases each. For these, we compute saliency maps (Wang et al. 2023) at each layer and measure the proportion of information flow from the key ICD to the generated answer tokens, relative to the total layer-wise flow. This yields the contribution score $s_{\text{contrib}}$. Comprehensive experimental details are provided in Appendix 2.
@@
Stage I operates on the shallow layer set $\mathcal{L}_{\text{stageI}}$, focusing on improving image-text alignment within each ICD to enhance LVLM’s early perception of key visual features. This not only mitigates the early attention deficits but also benefits subsequent layers. To achieve this goal, we first need to identify the key visual tokens in each ICD based on the semantics of its paired Q-A text. However, conventional metrics, such as summing attention scores from text to image tokens or relying on embedding similarity, are inadequate for multimodal ICL, as they may exacerbate attention deficits or fail to capture the nuanced semantics of Q-A dialogues (Marino et al. 2021). To address this, we introduce a dynamic attention increment strategy tailored to multimodal ICDs.
@@
$P_{l}(\mathbf{T}_{i})=\operatorname{softmax}_{j}\Bigl{(}\tfrac{1}{H}\sum_{h=1}^{H}A^{l,h}(\mathbf{T}_{i},j)\Bigr{)}\in\mathbb{R}^{1\times\mathbf{S}^{I}_{i}},$ (3)
@@
Next, we calculate the differences between these distributions as dynamic attention increments. They provide bias-reduced estimates of the image tokens’ contributions to both question understanding and answer making within each ICD. We quantify them by defining two non-negative forward gains $c^{1}_{l,i},c^{2}_{l,i}\in\mathbb{R}^{1\times\mathbf{S}^{I}_{i}}$in a divergence-like form:
@@
where $\cdot$$_{+}$ keeps only the positive part. After summing the two gains, column $j$ is the score of the $j$-th image token:
@@
$A^{l,h}(r,j)\leftarrow\ A^{l,h}(r,j)+\,\frac{n-i+1}{n}\,\frac{s_{i,j}}{\displaystyle\max_{j^{\prime}\in\mathcal{K_{I}}_{i}}s_{i,j^{\prime}}}.$ (7)
where $j\in\mathcal{K_{I}}_{i}$ and $r$ denotes any later token in $X$. The factor (ni+1)/n(n-i+1)/n is used to offset the position bias noted in Finding 3. Note that during this stage, we also detect and enhance the query sample’s key image tokens through only $c^{1}_{l,n-1}$. With intra‑ICD grounding, image tokens that align well with corresponding textual semantics receive greater attention in subsequent layers and during answer generation, providing a bias-corrected basis for deeper reasoning.
@@
After highlighting the most semantically aligned image tokens in Stage I, Stage II works on the middle layers $\mathcal{L}_{\text{stageII}}$ to refine the global information flow under the guidance of the query. Inspired by (Singh et al. 2024), which shows that specific attention heads dominate the information flow from the query sample to ICDs during ICL, we perform fine‑grained and targeted modulation at the head level to lessen the influence of other complex interactions in multimodal ICL.
We first use the attention flow from the query sample to the context (i.e., ICDs) as a signal to identify query-centric heads. As information gradually shifts toward later tokens in the middle layers, and text typically exhibits a stronger attention‑flow tendency (Gao et al. 2019), we measure the query→context flow using only the query sample’s text. Recall that $\mathbf{S}^{T}_{n+1}$ denotes the text token index set of the query sample. For each head $h$ in layer $l\!\in\!\mathcal{L}_{\text{stageII}}$, we compute:
$\rho^{l,h}=\frac{1}{|\mathbf{S}^{T}_{n+1}|}\sum_{q\in\mathbf{S}^{T}_{n+1}}\sum_{c\in\mathbf{S}_{ctx}}A^{l,h}(q,c),$ (8)
@@
For the embeddings of the ii‑th ICD taken from $\mathcal{L}_{\text{stageI}}[-1]$, we compute the mean of the key image tokens $\mathcal{K_{I}}_{i}$ to obtain a visual vector and the mean of all question‑and‑answer tokens to obtain a textual vector. We then concatenate these two vectors and apply $\ell_{2}$-normalization to the resulting embedding, producing the joint representation $p_{i}$. We apply the same procedure to the query sample to obtain $p_{\text{query}}$. The cosine similarity between $p_{i}$ and $p_{\text{query}}$ serves as a query-centric score:
@@
$A^{l,h}(r,J)\;\leftarrow\;A^{l,h}(r,J)+\frac{n-i+1}{n}\,w_{i}.$ (10)
We apply this modulation to all key image tokens and text tokens of each ICD to maintain completeness, so $J\in(\mathcal{K_{I}}_{i}\cup\mathbf{S}^{Q}_{i}\cup\mathbf{S}^{A}_{i})$ and $r$ denote any subsequent token in $X$. (ni+1)/n(n-i+1)/n applies the same position decay as in Eq.7. Stage II enhances the attention given to each ICD in proportion to its contribution, ensuring that key information is not lost within the extensive context. CAMA’s two stages jointly enable the LVLM to exploit the provided context more effectively. All remaining layers maintain their original attention.
@@
Benchmarks and models. Following the standard multimodal ICL evaluation (Awadalla et al. 2023), we test CAMA on VQAv2 (Goyal et al. 2017), VizWiz (Gurari et al. 2018), and OK-VQA (Marino et al. 2019). To further assess the generalization of CAMA, we also evaluate it on GQA (Hudson and Manning 2019), TextVQA (Singh et al. 2019), the CLEVR subset of VL-ICL bench (Zong, Bohdal, and Hospedales 2025), and MMStar (Chen et al. 2024a). In addition to LLaVA-NeXT-7B and Idefics2-8B, we also report results on two latest LVLMs, InternVL2.5-8B (Chen et al. 2025) and Qwen2.5VL-7B (Bai et al. 2025).
Baselines. We compare CAMA with five baselines. (1) Vanilla denotes the vanilla models. (2) The instruction‑augmented method (+Inst) add an instruction before each sequence: “First, study the examples we provide. Then utilize what you have learned to answer the new question.” (3) Contrastive decoding (CD) (Lee, Tsai, and Chiu 2024) replaces each ICD image with a blank one and uses the distorted logits to calibrate the original logits. (4) Visual enhancement (VE) (Su et al. 2025) manually draws a red bounding box around the relevant region of each ICD image. (5) SoFt Attention (SoFA) (Tian et al. 2025) is a training‑free method that inserts a bidirectional attention mask after every two decoder layers, which reduces position bias when multiple images are contained in the input.
@@
Implementation details. For each benchmark, samples in its validation set act as query samples, each paired with eight randomly retrieved ICDs from the training split, forming an 8-shot sequence. Stage I is applied to the 2nd and 3rd layers. Stage II is applied to every second layer from the 7th through the 19th. We set $k_{\text{I}}=k_{\text{II}}=20$. All experiments are conducted on 4 NVIDIA H200 GPUs.
@@
Adaptability to diverse sequence configurations. The in‑context sequence configuration is key to ICL performance. Our main experiments build sequences using uniform random sampling (RS), whereas some advanced systems leverage embedding similarity to further improve sequence quality. Each candidate ICD and the query sample are encoded using CLIP-ViT-L/14 to compute cosine similarity, and the top eight matches are selected as ICDs. I2I uses image embeddings only, while IQ2IQ concatenates image and question embeddings. We use TACO (Li et al. 2025c), a SOTA language-model-based ICD retriever. As shown in Figure 4(b), CAMA consistently improves accuracy by 2.59% to 5.08%. Stronger retrieval brings larger gains, indicating that CAMA’s activation effect also applies to configuration methods. This compatibility allows CAMA to integrate smoothly with future ICL advances.
@@
Case Study. In the case of Figure 1, we observe that CAMA reshapes attention, guiding the model to focus on visual features aligned with the text and relevant to the query. This shows the depth of CAMA’s ability to address attention deficits. We provide additional qualitative visualizations and failure case analysis in Appendix 3.4. For challenging tasks like CLEVR with vague text, Stage I’s dynamic attention increment helps locate key tokens. For MMStar, which involves complex long-text tasks, CAMA enables deeper reasoning. This highlights CAMA’s potential in areas like multimodal Chain-of-Thought (CoT) (Zhang et al. 2024).

### v3

# Make LVLMs Focus: Context-Aware Attention Modulation for Better Multimodal In-Context Learning

@@ Abstract
Multimodal in-context learning (ICL) is becoming a key capability that allows large vision-language models (LVLMs) to adapt to novel tasks without parameter updates, which expands their usefulness in many real-world applications. However, ICL performance remains unstable even when the in-context demonstrations (ICDs) are well matched, showing that LVLMs still struggle to make full use of the provided context. While existing work mainly focuses on prompt engineering or post-hoc logit calibration, we study the attention mechanisms inside LVLMs to address their inherent limitations. We identify two important weaknesses in their self-attention that hinder effective ICL. To address these weaknesses, we propose Context-Aware Modulated Attention (CAMA), a training-free and plug-and-play method that dynamically adjusts attention logits based on the input in-context sequence. CAMA uses a two-stage modulation process that strengthens attention to semantically important tokens, especially visual ones. Across four LVLMs and seven benchmarks, CAMA consistently outperforms vanilla models and baselines, showing clear effectiveness and generalization. It can also activate the intended benefits of prompt engineering methods and remains robust across different sequence configurations. Therefore, CAMA opens up new directions for improving multimodal reasoning through a deeper understanding of attention dynamics.
@@ Introduction
Large vision-language models (LVLMs) have emerged as powerful tools for multimodal information processing and generation (LVLM). Through large-scale pretraining, they integrate visual and textual signals into the shared representation space of large language models (LLMs) and have achieved notable success across vision-language tasks (lvlm1; lvlm2). However, adapting LVLMs to new domains remains challenging due to the high costs of multimodal data preparation and training.
To mitigate these costs, researchers are applying in-context learning (ICL), a technique widely used in LLMs (iclr1; iclr2), to LVLMs (icl3). In ICL, a few in-context demonstrations (ICDs) are incorporated into the input as reference examples, enabling the model to adapt to new tasks by interpreting these demonstrations without parameter updates. Recent advancements in model architectures and training protocols have enabled LVLMs to process multiple images and perform interleaved reasoning, making multimodal ICL practical (flamin; idefics; qwen; intern). These advances expand the practical scope of LVLMs (icl1; icl2; re1).
@@
However, the benefits of multimodal ICL can be diminished by its pronounced instability. Recent studies have found that LVLM ICL performance is highly sensitive to subtle prompt details. Minor changes in the order or formatting of ICDs can cause significant performance swings (sen1; sen2; sen3). This sensitivity intensifies as the number of ICDs grows or when complex multimodal reasoning is required (unstable; icv2). To improve multimodal ICL in LVLMs, two main routes have been explored. The first optimizes the prompt. It adds guiding text, highlights key image regions, or selects and orders ICDs with predefined metrics (conf1; conf2). However, prompt engineering requires substantial prior knowledge, lacks stability, and depends heavily on the base model. The second route edits models’ internal logits, as in contrastive decoding. This approach alters reasoning more directly but needs distorted inputs and extra forward passes for calibration (attn1; contrastive). Since practical ICL prioritizes efficiency, our work investigates the following questions: What is the intrinsic limitation underlying the instability of multimodal ICL in LVLMs? Can we address it using a more efficient, training-free method?
In pursuit of these questions, we examine the attention dynamics of LVLMs during multimodal ICL. Because their inputs interleave images and texts, we move beyond the existing focus on single-image scenarios and analyze LVLMs from two distinct perspectives: (1) text-based visual grounding within each image-text pair, and (2) query-sample-driven attention allocation across ICDs. We visualize attention deficits in both aspects through targeted experiments and verify that these deficits lead to insufficient utilization of the in-context sequence by the LVLM, resulting in unstable multimodal ICL. In response to these deficits, we present Context-Aware Modulated Attention (CAMA), a training-free method that dynamically modulates the model’s internal attention logits during inference based on the input context. CAMA addresses the two deficits with a two-stage modulation that targets the shallow and middle layers of the decoder. Stage I applies intra-ICD grounding, highlighting the image tokens in each ICD and in the query sample that best align with their accompanying text, which reduces the attention sink introduced by the interleaved image-text format. Stage II performs query-centric routing, reallocating attention among ICDs in proportion to their value to produce the desired answer, thus improving the use of context. Experiments demonstrate that CAMA delivers consistent performance gains for multimodal ICL, providing insights for reshaping attention mechanisms to build models with better visual capabilities.
@@
In-context Learning (ICL). ICL enables models to solve unseen tasks by conditioning on an input sequence of input–output examples (i.e., ICDs) without updating any parameters (iclr1; iclr2). This capability markedly improves their practicality and is especially valued in resource‑intensive multimodal domains (iclr4). Large vision–language models (LVLMs) gain ICL capabilities through targeted pretraining or fine‑tuning on interleaved image–text data, as in Flamingo (flamin) and LLaVA-NeXT (llava). Now, it has become a core ability of commercial model families such as QwenVL (qwen) and InternVL (intern). Consequently, recent work has begun to explore the mechanisms of multimodal ICL (icl1; icl2). However, few studies probe internal attention. Therefore, cross-modal interactions in multimodal ICL remain only partially understood, which prevents state-of-the-art (SOTA) models from fully exploiting ICL’s potential.
Enhancing Multimodal ICL. Limited use of input‑sequence information by LVLMs is considered a major source of instability in multimodal ICL (sen1; sen2; sen4). Three solution paths have thus emerged. First, task-oriented datasets combined with instruction tuning (tuning1; tuning2) or direct preference optimization (DPO) (DPO) train LVLMs to reason across multiple ICDs. Second, prompt-level methods optimize ICD selection with metrics such as similarity scores and information entropy (conf1; conf2; conf3; order1), or by training automatic selectors (taco; lever). Third, calibration-based methods adjust LVLMs’ final logits using contrastive decoding (cd1; cd2; cd3). While these approaches improve performance, they all face challenges. The first requires extensive curated data and parameter updates, the second is less adaptable to fixed-prompt scenarios, and both depend on the base model. The third needs to design distorted input sequences, plus multiple forward passes. In contrast, our CAMA sidesteps all these drawbacks by modulating attention logits at inference time without additional data or tuning.
@@
which directly reveals both the direction and intensity of interactions between any two tokens during the forward pass. After applying a causal mask followed by softmax, $\mathbf{A}^{l,h}$ becomes the attention weight matrix that is multiplied with the value $V^{l,h}$ to produce the head’s attention output.
@@
To understand the instability of multimodal in-context learning in LVLMs, we delve into the model’s attention mechanism to analyze its dynamics in both effective and ineffective scenarios. Our objective is to uncover distinct patterns that distinguish these cases. Specifically, we investigate attention dynamics at two levels: (1) Within each ICD: We examine text-based visual grounding to verify whether the model pays attention to relevant objects in the images. (2) Across ICDs: We evaluate whether the model effectively distributes attention among different ICDs according to their relevance to the query. To ensure generalizable experimental results, we experiment with 3-shot settings on two LVLMs: Llava-NeXT-7B (llava) and Idefics2-8B (idefics2). Query samples and ICDs are taken from the validation and training sets of VQAv2, respectively.
Setups. We pair each query with three ICDs of the same question type (e.g., “How many” or “Is there”) and process the resulting sequences using both LVLMs. From this pool, we identify 2,500 sequences where LLaVA-NeXT-7B produces a correct answer while Idefics2-8B fails, and another 2,500 sequences with the opposite outcome. These form two distinct sets of 5,000 cases each: a “correct” group (indicating effective ICL) and a “wrong” group (indicating ineffective ICL). For each sequence, we manually annotate bounding boxes on the ICD images based on their corresponding textual descriptions. We then extract attention maps from each model layer, identify the top 20% most-attended regions, and compute their Intersection over Union (IoU) with the annotated boxes to obtain the intra-ICD alignment score $s_{\text{align}}$. Additionally, we sample 10,000 sequences from the original pool and replace two of the three ICDs in each sequence with unrelated images, leaving only one key ICD relevant to the query type. For each of these sequences, we generate three variants by positioning the key ICD in the first, second, and third slots, respectively. Using the same processing pipeline, we construct new correct and wrong groups of 5,000 cases each. For these, we compute saliency maps (sail) at each layer and measure the proportion of information flow from the key ICD to the generated answer tokens, relative to the total layer-wise flow. This yields the contribution score $s_{\text{contrib}}$. Experimental details are provided in Appendix 2.
@@
Stage I operates on the shallow layer set $\mathcal{L}_{\text{stageI}}$, focusing on improving image-text alignment within each ICD to enhance LVLM’s early perception of key visual features. This not only mitigates the early attention deficits but also benefits subsequent layers. To achieve this goal, we first need to identify the key visual tokens in each ICD based on the semantics of its paired Q-A text. However, conventional metrics, such as summing attention scores from text to image tokens or relying on embedding similarity, are inadequate for multimodal ICL, as they may exacerbate attention deficits or fail to capture the nuanced semantics of Q-A dialogues (qa). To address this, we introduce a dynamic attention increment strategy tailored to multimodal ICDs.
@@
$P_{l}(\mathbf{T}_{i})=\operatorname{softmax}_{j}\Bigl(\tfrac{1}{H}\sum_{h=1}^{H}\textbf{A}^{l,h}(\mathbf{T}_{i},j)\Bigr)\in\mathbb{R}^{1\times S^{I}_{i}},$ (3)
@@
Next, we calculate the differences between these distributions as dynamic attention increments. They provide bias-reduced estimates of the image tokens’ contributions to both question understanding and answer making within each ICD. We quantify them by defining two non-negative forward gains $c^{1}_{l,i},c^{2}_{l,i}\in\mathbb{R}^{1\times S^{I}_{i}}$in a divergence-like form:
@@
where $\cdot$$_{+}$ keeps only the positive values by setting all negative ones to zero. After summing the two gains, column $j$ is the score of the $j$-th image token:
@@
$\textbf{A}^{l,h}(r,j)\leftarrow\ \textbf{A}^{l,h}(r,j)+\,\frac{n-i+1}{n}\,\frac{s_{i,j}}{\displaystyle\max_{j^{\prime}\in\mathcal{K_{I}}_{i}}s_{i,j^{\prime}}+\epsilon}.$ (7)
where $j\in\mathcal{K_{I}}_{i}$ and $r$ denotes any later token in $X$. The factor (ni+1)/n(n-i+1)/n is used to offset the position bias noted in Finding 3. Note that during this stage, we also detect and enhance the query sample’s key image tokens through only $c^{1}_{l,n+1}$. With intra‑ICD grounding, image tokens that align well with corresponding textual semantics receive greater attention in subsequent layers and during answer generation, providing a bias-corrected basis for deeper reasoning.
@@
After highlighting the most semantically aligned image tokens in Stage I, Stage II works on the middle layers $\mathcal{L}_{\text{stageII}}$ to refine the global information flow under the guidance of the query. Inspired by (head), which shows that specific attention heads dominate the information flow from the query sample to ICDs during ICL, we perform fine‑grained and targeted modulation at the head level to lessen the influence of other complex interactions in multimodal ICL.
We first use the attention flow from the query sample to the context (i.e., ICDs) as a signal to identify query-centric heads. As information gradually shifts toward later tokens in the middle layers, and text typically exhibits a stronger attention‑flow tendency (flow), we measure the query→context flow using only the query sample’s text. Recall that $\mathbf{S}^{T}_{n+1}$ denotes the text token index set of the query sample. For each head $h$ in layer $l\!\in\!\mathcal{L}_{\text{stageII}}$, we compute:
$\rho^{l,h}=\frac{1}{|\mathbf{S}^{T}_{n+1}|}\sum_{q\in\mathbf{S}^{T}_{n+1}}\sum_{c\in\mathbf{S}_{ctx}}\textbf{A}^{l,h}(q,c),$ (8)
@@
For the embeddings of the ii‑th ICD taken from $\mathcal{L}_{\text{stageI}}[-1]$, we compute the mean of the key image tokens $\mathcal{K_{I}}_{i}$ to obtain a visual vector and the mean of all question‑and‑answer tokens to obtain a textual vector. We then concatenate these two vectors and apply $\ell_{2}$-normalization to the resulting embedding, yielding the joint representation $p_{i}$. We apply the same steps to the query sample to obtain $p_{\text{query}}$. The cosine similarity between $p_{i}$ and $p_{\text{query}}$ is the query-centric score:
@@
$\textbf{A}^{l,h}(r,J)\;\leftarrow\;\textbf{A}^{l,h}(r,J)+\frac{n-i+1}{n}\,w_{i}.$ (10)
We apply this modulation to all key image tokens and text tokens of each ICD to maintain completeness, so $J\in(\mathcal{K_{I}}_{i}\cup\mathbf{S}^{Q}_{i}\cup\mathbf{S}^{A}_{i})$ and $r$ denote any subsequent token in $X$. (ni+1)/n(n-i+1)/n applies the same position decay as in Eq.7. Stage II enhances the attention given to each ICD in proportion to its contribution, ensuring that key information is not lost within the extensive context. The two stages of CAMA jointly enable the LVLM to exploit the provided context more effectively. All remaining layers maintain their original attention.
@@
Benchmarks and models. Following the standard multimodal ICL evaluation (openflamingo), we test CAMA on VQAv2 (vqav2), VizWiz (VizWiz), and OK-VQA (OK-VQA). To further assess the generalization of CAMA, we also evaluate it on GQA (GQA), TextVQA (textvqa), the CLEVR subset of VL-ICL bench (vlicl), and MMStar (mmstar). In addition to LLaVA-NeXT-7B and Idefics2-8B, we also report results on two latest LVLMs, InternVL2.5-8B (intern) and Qwen2.5VL-7B (qwen).
Baselines. We compare CAMA with five baselines. (1) Vanilla denotes the vanilla models. (2) The instruction‑augmented method (+Inst) add an instruction before each sequence: “First, study the examples we provide. Then utilize what you have learned to answer the new question.” (3) Contrastive decoding (CD) (contrastive) replaces each ICD image with a blank one and uses the distorted logits to calibrate the original logits. (4) Visual enhancement (VE) (enhance) manually draws a red bounding box around the relevant region of each ICD image. (5) SoFt Attention (SoFA) (sofa) is a training‑free method that inserts a bidirectional attention mask after every two decoder layers, which reduces position bias when multiple images are contained in the input.
@@
Implementation details. For each benchmark, samples in its validation set act as query samples, each paired with eight randomly retrieved ICDs from the training split, forming an 8-shot sequence. Stage I is applied to the 2nd and 3rd layers. Stage II is applied to every second layer from the 7th through the 19th. We set $k_{\text{I}}=k_{\text{II}}=20$. All experiments are conducted on NVIDIA H200 GPUs.
@@
Adaptability to diverse sequence configurations. The in‑context sequence configuration is key to ICL performance. Our main experiments build sequences using uniform random sampling (RS), whereas some advanced systems leverage embedding similarity to improve sequence quality further. Each candidate ICD and the query sample are encoded using CLIP-ViT-L/14 to compute cosine similarity, and the top eight matches are selected as ICDs. I2I uses image embeddings only, while IQ2IQ concatenates image and question embeddings. We also adopt TACO (taco), a SOTA language-model-based ICD retriever. As shown in Figure 4(b), CAMA consistently improves accuracy by 2.59% to 5.08%. Stronger retrieval brings larger gains, indicating that CAMA’s activation effect also applies to configuration methods. This compatibility allows CAMA to integrate smoothly with future ICL advances.
@@
Case Study. In the case of Figure 1, we observe that CAMA reshapes attention, guiding the model to focus on visual features aligned with the text and relevant to the query. This shows the depth of CAMA’s ability to address attention deficits. We provide additional qualitative visualizations and failure case analysis in Appendix 3.4. For challenging tasks like CLEVR with vague text, Stage I’s dynamic attention increment helps locate key tokens. For MMStar, which involves complex long-text tasks, CAMA enables deeper reasoning. This highlights CAMA’s potential in complicated or reasoning-intensive scenarios, such as multimodal Chain-of-Thought (CoT) (cot).